OpenStudy (anonymous):

Q is the point of intersection of the diagonals of one face of a cube with edges 2 units. The length (in units) of QR is _____.

OpenStudy (anonymous):



OpenStudy (anonymous):

@aum

OpenStudy (aum):

QR^2 = RS^2 + SQ^2; RS = 2 SQ = 1/2 SP SP^2 = 2^2 + 2^2

OpenStudy (anonymous):

8

OpenStudy (anonymous):

why sp^2=2^2 + 2^2?

OpenStudy (aum):

SP is the diagonal of a square with sides 2 and 2. Apply Pythagoras Theorem.

OpenStudy (anonymous):

only =2^2

OpenStudy (aum):